OpenStudy (anonymous):

how do you integrate tan^6(x)dx

OpenStudy (anonymous):

\[\int\limits \tan ^{6}dx\]

OpenStudy (turingtest):

\[\tan^6x=\tan^2x\cdot\tan^4x=\tan^2x(\sec^2x-1)^2\]expand and u-sub

OpenStudy (nenadmatematika):

Go TuringTest!!!!! :D

OpenStudy (turingtest):

\[\tan^6x=\tan^4x\cdot\tan^2x=\tan^4x(\sec^2x-1)=\tan^4x\sec^2x-\tan^4x\]\[=\tan^4x\sec^2x-\tan^2(\sec^2-1)=\tan^4x\sec^2x-\tan^2\sec^2+\tan^2x\]\[=\tan^4x\sec^2x-\tan^2\sec^2+\sec^2x-1\]\[u=\tan x\to du=\sec^2xdx\]I bet there is a faster way though...

OpenStudy (turingtest):

@UTLEO got it? are we good?

OpenStudy (anonymous):

I am trying to understand it =]

OpenStudy (nenadmatematika):

UTLEO, be careful, he wants a medal...:D haha....

OpenStudy (anonymous):

LOL

OpenStudy (turingtest):

say it was\[\int\tan^4xdx\]I'm utilizing the identity\[\tan^2x=1-\sin^2x\]watch what happens:\[\int\tan^4xdx=\int\tan^2x\cdot\tan^2xdx=\int\tan^2x(\sec^2x-1)dx\]\[=\int\tan^2x\sec^2xdx-\int\tan^2xdx\]now the first integral can be done with a u-sub\[u=\tan x\to du=\sec^2xdx\]so the first integral becomes\[\int\tan^2x\sec^2xdx=\int u^2du=\frac13\tan^3x+C\]right?

OpenStudy (anonymous):

Great answer!!

OpenStudy (turingtest):

for the other integral we can use the same identity\[\int\tan^2xdx=\int\sec^2x-1dx=\tan x-x+C\] so the whole thing is\[\int\tan^4xdx= \frac13\tan^3x-\tan x+x+C\]right? now look at what I did above for tan^6 I used the exact same procedure, but had to repeat it one extra time..

OpenStudy (turingtest):

glad to help, see you around :D

OpenStudy (anonymous):

I have one more "tricky" question to me please help me for this one \[\int\limits \cos pixcos(\pi x/2)dx\]

OpenStudy (turingtest):

\[\int\cos(\pi x)\cos(\frac{\pi x}2)dx\]?

OpenStudy (anonymous):

yea

OpenStudy (turingtest):

there is a simple trig identity that fixes so many of these types of integrals...

OpenStudy (turingtest):

\[\cos\alpha\cos\beta=\frac12[\cos(\alpha-\beta)+\cos(\alpha+\beta)]\]now the integral should be no problem ;)

OpenStudy (anonymous):

haven't thought of that, thanks for reminding!

OpenStudy (turingtest):

those sina*cosb sina*sinb cosa*cosb ones always get people don't be one of them, remember those identities!

OpenStudy (anonymous):

definitely!